OpenStudy (samara8954):

Sandy works at a clothing store. She makes $9 per hour plus earns 15% commission on her sales. She worked 80 hours over the last two weeks and had a total of $2,744 in sales before taxes. Which of the following is closest to how much she will earn in hourly wages and commission for those two weeks? $520 $720 $412 $1,132 @mathmate @dtan5457

OpenStudy (samara8954):

@mathmate

OpenStudy (anonymous):

which piece of equipment should the student use to measure the water he gives to each plant

OpenStudy (anonymous):

plz hlp

OpenStudy (samara8954):

does anyone know the answer

OpenStudy (samara8954):

help

OpenStudy (samara8954):

I will medal

OpenStudy (anonymous):

it is 412

OpenStudy (anonymous):

I just got it guys

OpenStudy (samara8954):

412 is wrong

OpenStudy (anonymous):

You know why it can't be anything lower than 720 right?

OpenStudy (samara8954):

maybe?

OpenStudy (anonymous):

If she makes $9 and hour, and she worked 80 hours in those two weeks. 9 * 80 = 720 So 720 is the lowest she could earn.

OpenStudy (anonymous):

Since she made $2,744 in sales and she earns 15% commission on her sales, you'll have to calculate 15% out of 2744.

OpenStudy (mathmate):

You already know that salary equals 80*$9=$720 You need to calculate the amount of commission she receives, which is \(2744*15\%=$(2744*\frac{15}{100})=?\) Add the commission to the base salary will give you what she should earn in those two weeks.